Find the transition matrix P and the diagonal matrix D such that A = P-DP, if [2 1 1] A 1 2 1 1 2
Answer #1

Az 2 The charadesstic car is b&+91-40 C&D) Cha=0 ~ 1=111,4 op eigenvalues Eigenvertas fatal (A-I) X=0 167-69 OxfXNXZ Optix a=)=28,7X27);=0 x^2X27X3 =0 X772-23 20 0-0 ->) -3Xp+3Y3 =0 @3) -3%,+3850 x=3 ** X ta x = ха ts -63 [i] Xz = is epenwets f3 1=